# Starobinsky Inflation with T-Model Kähler Geometries

### v1

@@ Abstract
We present novel implementations of Starobisky-like inflation within Supergravity adopting Kähler potentials for the inflaton which parameterize hyperbolic geometries known from the T-model inflation. The associated superpotentials are consistent with an $R$ and a global or gauge $U(1)_{X}$ symmetries. The inflaton is represented by a gauge singlet or non-singlet superfield and is accompanied by a gauge-singlet superfield successfully stabilized thanks to its compact contribution into the total Kähler potential. Keeping the Kähler manifold intact, a conveniently violated shift symmetry is introduced which allows for a slight variation of the predictions of Starobinsky inflation: The (scalar) spectral index exhibits an upper bound which lies close to its central observational value whereas the constant scalar curvature of the inflaton-sector Kähler manifold increases with the tensor-to-scalar ratio.
@@ 1 Introduction
For $n=n_{\rm d}=2$, $V_{\rm I}$ has been motivated in Ref. [29, 60] via a breaking of conformal symmetry in a non-SUSY framework. It also appears in unified models of no-scale $\alpha$-SI with SUSY breaking [65]. Here we extend the analysis in the context of SUGRA providing a method which allows the generation of $V_{\rm I}$ from conveniently selected $W$ and $K$. Our particle content includes besides the inflaton superfield(s) an extra gauge singlet superfield which assist in the stabilization of the SUGRA potential during SI [38]. Moreover, we employ monomial $W$’s consistent with an $R$ and a global or gauge $U(1)_{X}$ symmetry. On the other hand, the $K$’s respect the $R$ and the gauge symmetry and are holomorphically equivalent to those yielding Eq. (2) [56, 66, 67, 68]. In other words, $K$ has the well-known form employed in T-model inflation [56, 66] up to a number of extra holomorphic and anti-holomorphic terms which do not influence the resulting Kähler metric. The one pair of these terms endows $K$ with a shift symmetry which facilitates the performance of inflation – cf. Ref. [56, 66, 67, 68]. The second pair provides a breaking of the aforementioned shift symmetry which is already violated mildly in $W$. Both violations are natural in the ’t Hooft sense [69] for low enough values of the exponent $n_{\rm d}$ in $K$ and the coupling constant $\lambda$ entered in $W$. The particular importance of an enhanced shift symmetry in taming the so-called $\eta$-problem of inflation in SUGRA is already recognized for gauge singlets, e.g., in Ref. [70, 74, 75, 73, 71, 72] and non-singlets, e.g., in Ref. [76, 77, 78, 79, 48, 36].
@@
Compared to the aforementioned implementations of SI within SUGRA, the present version assures a totally symmetric Kähler manifold together with a simple $W$. Let us recall that polynomial $W$’s with $R$ symmetry are reconciled with a $K$ which does not parameterize specific Kähler manifold in Ref. [58] whereas totally symmetric $K$’s usually require complicate $W$’s – see e.g. Ref. [20, 56, 57, 61, 30]. On the contrary, the employed here $W$’s are very common in particle physics and so the inflaton can be easily identified with a field already present in the theory, e.g, the right-handed sneutrino [70, 80, 81] or a superheavy Higgs superfield responsible for the spontaneous breaking of a gauge symmetry [36, 37]. Also, it is expected that this scheme assures naturally a low enough reheating temperature, potentially consistent with the gravitino constraint and non-thermal leptogenesis [31, 36, 37, 83, 84] if connected with a version of the Minimal SUSY Standard Model (MSSM). Furthermore, our proposal does not require tuning of parameters – besides a mild one in the initial conditions –, it is not based on any conjecture such as that of induced gravity [31, 32, 33, 36, 37, 34, 10, 35] and provides a relative flexibility as regards the observables. At last, it offers us the opportunity to exemplify the Kähler potential engineering which allows to obtain the various desired factors of $V_{\rm I}$ in Eq. (3) together with a desirable Kähler metric.
@@
We below describe how we can formulate SI in the context of SUGRA in Sec. 2 and we specify two versions of SI: one employing a gauge-singlet inflaton in Sec. 3 called for short Chaotic Starobinsky Inflation (CSI) and one with a gauge non-singlet inflaton leading to the breaking of a gauge group called Higgs Starobinsky Inflation (HSI) in Sec. 4. Our conclusions are summarized in Sec. 6. In Appendix 6 we demonstrate that our proposed $K$’s enjoy an enhanced shift symmetry. Throughout the text, the subscript $,\chi$ denotes derivation with respect to (w.r.t) the field $\chi$ and charge conjugation is denoted by a star (). Unless otherwise stated, we use units where the reduced Planck scale $m_{\rm P}=2.4\cdot 10^{18}~{}{\mbox{\rm GeV}}$ is set equal to unity.
@@ 2.1 General Set-up
where $\mathcal{R}$ is the space-time Ricci scalar curvature, $\mathfrak{g}$ is the determinant of the Friedmann-Robertson-Walker metric, $g_{\mu\nu}$, with signature (+,,,)(+,-,-,-)( + , - , - , - ). Also, summation is taken over the scalar fields $z^{\alpha}$, the kinetic mixing of which is controlled by the Kähler potential $K$ and the relevant metric defined as
@@
with $A^{\rm a}_{\mu}$ being the vector gauge fields, $g$ the (unified) gauge coupling constant and $T^{\rm a}$ with ${\rm a}=1,...,\mbox{\sf\footnotesize dim}G_{\rm GUT}$ the generators of a gauge group $G_{\rm GUT}$. Here and henceforth, the scalar components of the various superfields are denoted by the same superfield symbol.
@@ 3.1 Set-up
$\mathcal{M}_{211}=(SU(2)/U(1))_{S}\times\left(SU(1,1)/U(1)\right)_{\Phi}\>\>\>\mbox{and}\>\>\>{\sf R}_{211}=-{2}/{N}+2/N_{S}.$ (18)
@@ 4.3 Inflationary Potential
Table 2: Mass spectrum for HSI along the inflationary trajectory of Eq. (23).
@@ 5.1 Analytic Results
In practice, for $n_{\rm d}<n/2$, $\phi_{1\rm f}<\phi_{2\rm f}$ and so SI terminates at $\phi_{\rm f}=\phi_{2\rm f}$ whereas for larger $n_{\rm d}$ the inverse hierarchy turns out and so $\phi_{\rm f}=\phi_{1\rm f}$.
@@
where $\delta n=n_{\rm d}-2n<0$. The last inequality stems from the fact that the dominant contribution to $I_{N}$ originates from the non-logarithmic term and $\phi<1$ due to the presence of a pole in $J$ – see Eq. (19) – and the effective nature of SUGRA. Consequently, the positivity of $N_{\star}$ implies the upper bound above on $\delta n$ which restricts seriously the allowed region of our model as we see in Sec. 5.2.
@@ 5.2 Numerical Results
From Fig. 3b we remark that for both $n_{\rm d}$ values the magnitudes of the two plateaus are of the order $10^{-10}$ which are similar to that obtained in pure SI [31, 32, 10, 36, 48, 49, 37, 58]. However, these are one order of lower than that obtained in Refs. [33], where $r$ is a little more enhanced. Indeed, as verified from the values listed in Table 3, the level of the inflationary plateau increases with $r$. Moreover, the inflationary scale $V_{\rm I}^{1/4}$ turns out to be well below $m_{\rm P}$ and so the semi-classical approximation, used in our analysis, is perfectly valid. Note that here $m_{\rm P}$ is undoubtedly the ultraviolet cut-off scale of the theory thanks to the absence of large coefficients in the $K$’s. Recall that such large coefficients are used in models of induced-gravity [31, 32, 35, 10, 36, 37] or non-minimal [48, 49] inflation and the aforementioned scale has to be determined after an expansion of ${\cal A}$ in Eq. (4a) about $\langle{\phi}\rangle$.
@@ 6 Conclusions
We presented a novel implementation of SI (i.e. Starobinsky inflation) in the context of SUGRA confining ourselves to models displaying a kinetic mixing in the inflaton sector with a pole of order two – see Eq. (2) – and a scalar potential shown in Eq. (3). We considered two classes of models (CSI and HSI) depending on whether the inflaton is included into a gauge singlet or two gauge non-singlet fields. CSI and HSI are relied on the superpotentials in Eqs. (14) and (24) respectively which respect an $R$ symmetry and include an inflaton-accompanying field which facilitates the establishment of SI. On the other hand, the Kähler potential’s respect the $R$ and gauge symmetries and parameterize hyperbolic internal geometries. Namely, $K$ for CSI is given in Eq. (16) whereas for HSI we considered two distinct $K$’s shown in Eq. (28). The Higgflaton in the last case implements the breaking of a gauge $U(1)_{X}$ symmetry at a scale which may assume a value compatible with the MSSM unification. All the models excellently match with the observations by restricting the free parameters to reasonably ample regions of values. In particular, the predicted $r$’s increase with $N$, while $n_{\rm s}$ lies close to its central observational value – see Eq. (54a) for CSI and Eqs. (55a) and (56a) for HSI. The present data on $\delta_{21}$ and the self-consistency of the models allows us to delineate the overall allowed regions for selected $n$ – see Fig. 2 and 5. Hopefully, a more accurate determination of $n_{\rm s}$ and $r$ by future experiments [94, 96, 95, 97] will assist us to single out the most favorable one from the proposed models.
@@
The central message of our work is that SI is not exclusively implemented by the E-model kinetic mixing in Eq. (1). It is also attainable via T-model normalization in Eq. (2) if it is considered in conjunction with the potential in Eq. (3). The method applied for the construction of our models can be extended to other SUGRA models as those motivated by D branes [98] or those which assist to obtain de Sitter vacua in line with current LHC results on SUSY as in Ref. [57, 65]. Moreover, it can employed for supersymmetrizing models obtained adopting the Palatini approach to gravity [99, 100, 101]. At last, we expect that our models of HSI admit a post-inflationary completion along the lines of Ref. [31, 37, 36, 48, 68] since the mass parameters in those models are similar to those found in Eqs. (55b) and (56b).

### v2

@@ Abstract
We present novel implementations of Starobisky-like inflation within Supergravity adopting Kähler potentials for the inflaton which parameterize hyperbolic geometries known from the T-model inflation. The associated superpotentials are consistent with an $R$ and a global or gauge $U(1)_{X}$ symmetries. The inflaton is represented by a gauge-singlet or non-singlet superfield and is accompanied by a gauge-singlet superfield successfully stabilized thanks to its compact contribution into the total Kähler potential. Keeping the Kähler manifold intact, a conveniently violated shift symmetry is introduced which allows for a slight variation of the predictions of Starobinsky inflation: The (scalar) spectral index exhibits an upper bound which lies close to its central observational value whereas the constant scalar curvature of the inflaton-sector Kähler manifold increases with the tensor-to-scalar ratio.
@@
Published in Universe 11, no. 3, 75 (2025)
Special Issue
Particle Physics and Cosmology: A Themed Issue in Honor of Professor Dimitri Nanopoulos
@@ 1 Introduction
For $n=n_{\rm d}=2$, $V_{\rm I}$ has been motivated in Ref. [29, 60] via a breaking of conformal symmetry in a non-SUSY framework. It also appears in unified models of no-scale $\alpha$-SI with SUSY breaking [65]. Here we extend the analysis in the context of SUGRA providing a method which allows the generation of $V_{\rm I}$ from conveniently selected $W$ and $K$. Our particle content includes, besides the inflaton superfield(s), an extra gauge-singlet superfield which assist in the stabilization of the SUGRA potential during SI [38]. Moreover, we employ monomial $W$’s consistent with an $R$ and a global or gauge $U(1)_{X}$ symmetry. On the other hand, the $K$’s respect the $R$ and the gauge symmetry and are holomorphically equivalent to those yielding Eq. (2) [56, 66, 67, 68]. In other words, $K$ has the well-known form employed in T-model inflation [56, 66] up to a number of extra holomorphic and anti-holomorphic terms which do not influence the resulting Kähler metric. The one pair of these terms endows $K$ with a shift symmetry which facilitates the performance of inflation – cf. Ref. [56, 66, 67, 68]. The second pair provides a breaking of the aforementioned shift symmetry which is already violated mildly in $W$. Both violations are natural in the ’t Hooft sense [69] for low enough values of the exponent $n_{\rm d}$ in $K$ and the coupling constant $\lambda$ entered in $W$. The particular importance of an enhanced shift symmetry in taming the so-called $\eta$-problem of inflation in SUGRA is already recognized for gauge singlets, e.g., in Ref. [70, 74, 75, 73, 71, 72] and non-singlets, e.g., in Ref. [76, 77, 78, 79, 48, 36].
@@
Compared to the aforementioned implementations of SI within SUGRA, the present version assures a totally symmetric Kähler manifold together with a simple $W$. Let us recall that polynomial $W$’s with $R$ symmetry are reconciled with a $K$ which does not parameterize specific Kähler manifold in Ref. [58] whereas totally symmetric $K$’s usually require complicate $W$’s – see e.g. Ref. [20, 56, 57, 61, 30]. On the contrary, the employed here $W$’s are very common in particle physics and so the inflaton can be easily identified with a field already present in the theory, e.g, the right-handed sneutrino [70, 80, 81, 82] or a superheavy Higgs superfield responsible for the spontaneous breaking of a gauge symmetry [36, 37]. Also, it is expected that this scheme assures naturally a low enough reheating temperature, potentially consistent with the gravitino constraint and non-thermal leptogenesis [31, 36, 37, 83, 84] if connected with a version of the Minimal SUSY Standard Model (MSSM). Furthermore, our proposal does not require tuning of parameters – besides a mild one in the initial conditions –, it is not based on any conjecture such as that of induced gravity [31, 32, 33, 36, 37, 34, 10, 35] and provides a relative flexibility as regards the observables. At last, it offers us the opportunity to exemplify the Kähler potential engineering which allows to obtain the various desired factors of $V_{\rm I}$ in Eq. (3) together with a desirable Kähler metric.
@@
We below describe how we can formulate SI in the context of SUGRA in Sec. 2 and we specify two versions of SI: one employing a gauge-singlet inflaton in Sec. 3 called for short Chaotic Starobinsky Inflation (CSI) and one with a gauge-non-singlet inflaton called Higgs Starobinsky Inflation (HSI) since it causes the breaking of a gauge symmetry – see Sec. 4. Our conclusions are summarized in Sec. 6. In Appendix 6 we demonstrate that our proposed $K$’s enjoy an enhanced shift symmetry. Throughout the text, the subscript $,\chi$ denotes derivation with respect to (w.r.t) the field $\chi$ and charge conjugation is denoted by a star (). Unless otherwise stated, we use units where the reduced Planck scale $m_{\rm P}=2.4\cdot 10^{18}~{}{\mbox{\rm GeV}}$ is set equal to unity.
@@ 2.1 General Set-up
where $\mathcal{R}$ is the space-time Ricci scalar curvature, $\mathfrak{g}$ is the determinant of the Friedmann-Robertson-Walker metric, $g_{\mu\nu}$, with signature (+,,,)(+,-,-,-)( + , - , - , - ). Also, summation is taken over the scalar fields $z^{\alpha}$ which are denoted by the same symbol of the corresponding superfield. The kinetic mixing of $z^{\alpha}$ is controlled by the Kähler potential $K$ and the relevant metric defined as
@@
with $A^{\rm a}_{\mu}$ being the vector gauge fields, $g$ the (unified) gauge coupling constant and $T^{\rm a}$ with ${\rm a}=1,...,\mbox{\sf\footnotesize dim}G_{\rm GUT}$ the generators of a gauge group $G_{\rm GUT}$.
@@ 3.1 Set-up
$\mathcal{M}_{211}=(SU(2)/U(1))_{S}\times\left(SU(1,1)/U(1)\right)_{\Phi}\>\>\>\mbox{and}\>\>\>{\sf R}_{211}=2/N_{S}-{2}/{N}.$ (18)
@@ 4.3 Inflationary Potential
Table 2: Mass spectrum for HSI with $K={\widetilde{K}_{2(11)^{2}\rm d}}$ and ${\widetilde{K}_{221\rm d}}$ along the inflationary trajectory of Eq. (23).
@@ 5.1 Analytic Results
In practice, for $n_{\rm d}<n/2$, $\phi_{1\rm f}<\phi_{2\rm f}$ and so SI terminates at $\phi_{\rm f}=\phi_{2\rm f}$ whereas larger $n_{\rm d}$’s give rise to the inverse hierarchy and so $\phi_{\rm f}=\phi_{1\rm f}$.
@@
where $\delta n=n_{\rm d}-2n<0$. The last inequality stems from the fact that the dominant contribution to $I_{N}$ originates from the non-logarithmic term with $\phi<1$. The presence of a pole in $J$ – see Eq. (19) – and the effective nature of SUGRA forces us to work in this domain of $\phi$. Consequently, the positivity of $N_{\star}$ implies the upper bound above on $\delta n$ which restricts seriously the allowed region of our model as we see in Sec. 5.2.
@@ 5.2 Numerical Results
From Fig. 3-(b) we remark that for both $n_{\rm d}$ values the magnitudes of the two plateaus are of the order $10^{-10}$ which are similar to that obtained in pure SI [31, 32, 10, 36, 48, 49, 37, 58]. However, these are one order of lower than that obtained in Refs. [33], where $r$ is a little more enhanced. Indeed, as verified from the values listed in Table 3, the level of the inflationary plateau increases with $r$. Moreover, the inflationary scale $V_{\rm I}^{1/4}$ turns out to be well below $m_{\rm P}$ and so the semi-classical approximation, used in our analysis, is perfectly valid. Note that here $m_{\rm P}$ is undoubtedly the ultraviolet cut-off scale of the theory thanks to the absence of large coefficients in the $K$’s. Recall that such large coefficients are used in models of induced-gravity [31, 32, 35, 10, 36, 37] or non-minimal [48, 49] inflation and the aforementioned scale has to be determined after an expansion of ${\cal A}$ in Eq. (4a) about $\langle{\phi}\rangle$.
@@ 6 Conclusions
We presented a novel implementation of SI (i.e. Starobinsky inflation) in the context of SUGRA confining ourselves to models displaying a scalar potential shown in Eq. (3) and a kinetic mixing in the inflaton sector with a pole of order two – see Eq. (2). We considered two classes of models – CSI, i.e., chaotic SI and HSI, i.e. Higgs SI – depending on whether the inflaton is included into a gauge-singlet or two gauge-non-singlet fields. CSI and HSI are relied on the superpotentials in Eqs. (14) and (24) respectively which respect an $R$ symmetry and include an inflaton-accompanying field which facilitates the establishment of SI. On the other hand, the Kähler potential’s respect the $R$ and gauge symmetries and parameterize hyperbolic internal geometries met in T-model inflationary settings. Namely, $K$ for CSI is given in Eq. (16) whereas for HSI we considered the two distinct $K$’s shown in Eq. (28). The Higgsflaton in the last case implements the breaking of a gauge $U(1)_{X}$ symmetry at a scale which may assume a value compatible with the MSSM unification. All the models excellently match with the observations by restricting the free parameters to reasonably ample regions of values. In particular, $n_{\rm s}$ lies close to its central observational value while $r$ increases with $N$ – see Eq. (54a) for CSI and Eqs. (55a) and (56a) for HSI. The present data on $\delta_{21}$ and the self-consistency of the models allows us to delineate the overall allowed regions for selected values of the exponent $n$ in Eq. (3) – see Fig. 2 and 5. The resulting allowed margin of $n_{\rm s}$ is extended from $0.961$ up to $0.969$, depending on the chosen $n$, and therefore it is less restrictive than the corresponding prediction of the traditional Starobinsky model in Eq. (48) – i.e. $n_{\rm s}\simeq 0.963$. Hopefully, a more accurate determination of $n_{\rm s}$ and $r$ by future experiments [94, 96, 95, 97] will assist us to single out the most favorable one from the proposed models.
@@
The central message of our work is that SI is not exclusively implemented by the E-model kinetic mixing in Eq. (1). It is also attainable via T-model normalization in Eq. (2) if it is considered in conjunction with the potential in Eq. (3). The method applied for the construction of our models can be extended to other SUGRA models as those motivated by D branes [98] or those which assist to obtain de Sitter vacua in line with current LHC results on SUSY as in Ref. [57, 65]. Moreover, it can be employed for supersymmetrizing models obtained adopting the Palatini approach to gravity [99, 100, 101].
We expect that our models of HSI admit a post-inflationary completion along the lines of Ref. [31, 37, 36, 48, 68] since the mass parameters in those models are similar to those found in Eqs. (55b) and (56b). As regards the ultraviolet completion, it would be interesting to investigate if our models belong to the string landscape or swampland [102]. Note that the swampland string conjectures are generically not satisfied in SUGRA-based models [103, 104] but there are suggestions [105, 106, 107] which may work in our framework too.